Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

National Small-bore Rifle Association

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject:  18th February 2002

A Chara,

I attach the entry forms for the Summer League Air Rifle postal competitions. I realise these are delayed but am asking that they be accepted as they were delayed by illness on my part.

Misé le meas,

Mark Dennehy, DURC Secretary